OpenStudy (anonymous):

A C1 = 6.0 µF and a C2 = 3.0 µF capacitor are connected in series, and this combination is connected in parallel with a C3 = 3.5 µF capacitor (see Fig. 19-52). What is the net capacitance?

OpenStudy (shane_b):

Using the formulas for series and parallel capacitance. Your final equation should look like thi: \[(\frac{1}{6}+\frac {1}{3})^{-1}+3.5=?\]

OpenStudy (anonymous):

5.5?

OpenStudy (shane_b):

Yes

OpenStudy (anonymous):

1/c = 1/c1 + 1/c2 + c3

OpenStudy (anonymous):

Thank you both :)

OpenStudy (shane_b):

np :)

OpenStudy (anonymous):

can i ask another question?

OpenStudy (anonymous):

yep